OpenStudy (anonymous):

Algebra 2 help? MEDAL + FAN Rationalize the denominator and simplify. 12/√7+2

OpenStudy (anonymous):

multiply the system by the complex conjugate \[\sqrt{7}-2/(\sqrt{7}-2)\]

OpenStudy (anonymous):

multiply nr and dnr with sqrt(7)-2

OpenStudy (anonymous):

\[\frac{ 12 }{ \sqrt{7}+2 }\frac{ \sqrt{7}-2 }{ \sqrt{7}-2 }\]

OpenStudy (kirbykirby):

\[ \frac{12}{\sqrt{7}+2}=\frac{12}{\sqrt{7}+2}\frac{\sqrt{7}-2}{\sqrt{7}-2}=\frac{12\sqrt{7}-24}{7-4}=4\sqrt{7}-8\]

OpenStudy (anonymous):

and it is not called the complex conjugate. Just the conjugate in this case

OpenStudy (anonymous):

Okay, I think I see it, but can you help me with another just so I can confirm that I understand how to do it? @kirbykirby

OpenStudy (kirbykirby):

ok sure

OpenStudy (anonymous):

6/√2-√3

OpenStudy (kirbykirby):

So this is similar, you have to multiply the top and bottom by the conjugate. The definition of the conjugate is: If you have \(a + b\), the conjugate is \(a - b\) If you have \(a-b\), then conjugate is \(a+b\) (you just change the middle sign) So, \[ \frac{6}{\sqrt{2}-\sqrt{3}}=\frac{6}{\sqrt{2}-\sqrt{3}}\frac{\sqrt{2}+\sqrt{3}}{\sqrt{2}+\sqrt{3}}\]

OpenStudy (anonymous):

Okay so that then turns to 6√2+6√3/√4-√3?

OpenStudy (kirbykirby):

the numerator is correct, but the denominator isn't

OpenStudy (anonymous):

Oops. I meant √4-√6. Is that right? Or would it be √4-√9?

OpenStudy (kirbykirby):

\(-\sqrt{9}\) since: \(\sqrt{3}\sqrt{3}=\sqrt{3\times 3}=\sqrt{9}\)

OpenStudy (anonymous):

Alright. And then after that is where I get stuck because I don't understand how to simplify it.

OpenStudy (kirbykirby):

so you have in the denominator: \(\sqrt{4}-\sqrt{9}=?\) can you take the square roots of these numbers?

OpenStudy (anonymous):

Would I just subtract 9 from 4?

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

\(\sqrt{4}=2\) \(\sqrt{9}=3\) Since \(2^2=4\) and \(3^2=9\)

OpenStudy (anonymous):

Ohhh! I see now.

OpenStudy (anonymous):

What do I do next? @@kirbykirby

OpenStudy (kirbykirby):

so in the denominator, 2 - 3 = -1 soo now you have \[ \frac{6\sqrt{2}+6\sqrt{3}}{-1}\] Now, recall: \[\frac{x+y}{a}=\frac{x}{a}+\frac{y}{a}\] so, you can separate your fraction as: \[ \frac{6\sqrt{2}}{-1}+\frac{6\sqrt{3}}{-1}\]